1. A population consists of six values (6, 9, 12, 15, 18, and 21). (with illustration)

e. Determine the mean, variance and standard deviation of the sample mean.



2. The scores of individual students on a national test have a normal distribution with mean of 18.6 and a standard deviation of 5.9. At Federico Ramos Rural High School, 76 students took the test. If the scores at this school have the same distribution as national scores, solve for the following: (with illustration)

a. determine the mean and standard deviation of the sampling distribution of the sample mean.


1
Expert's answer
2022-05-16T23:41:08-0400

1. We have population values 6,9,12,15,18,21, population size N=6 and sample size n=3.

Mean of population "(\mu)" = "\dfrac{6+9+12+15+18+21}{6}=13.5"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{1}{6}(56.25+20.25+2.25"


"+2.25+20.25+56.25)=26.25"

"\sigma=\sqrt{\sigma^2}=\sqrt{26.25}\approx5.1235"

Select a random sample of size 3 without replacement. We have a sample distribution of sample mean.

The number of possible samples which can be drawn without replacement is "^{N}C_n=^{6}C_3=20."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 6,9,12 & 9 \\
 \hdashline
 2 & 6,9,15 & 10 \\
 \hdashline
 3 & 6,9,18 & 11 \\
 \hdashline
 4 & 6,9,21 & 12 \\
 \hdashline
 5 & 6,12,15 & 11 \\
 \hdashline
 6 & 6,12,18 & 12 \\
 \hdashline
 7 & 6,12, 21 & 13 \\
 \hdashline
 8 & 6,15,18 & 13 \\
 \hdashline
 9 & 6,15,21 & 14 \\
 \hdashline
 10 & 6, 18,21 & 15 \\
 \hdashline
 11 & 9,12,15 & 12 \\
 \hdashline
 12 & 9, 12,18 & 13 \\
 \hdashline
 13 & 9, 12, 21 & 14 \\
 \hdashline
 14 & 9,15,18 & 14 \\
 \hdashline
 15 & 9,15,21 & 15 \\
 \hdashline
 16 & 9, 18,21 & 16 \\
 \hdashline
 17 & 12, 15,18 & 15 \\
 \hdashline
 18 & 12, 15,21 & 16 \\
 \hdashline
 19 & 12, 18,21 & 17 \\
 \hdashline
 20 & 15, 18,21 & 18 \\
 \hdashline
\end{array}"



B.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) & \bar{X}^2f(\bar{X})
\\ \hline
 9 & 1/20 & 9/20 & 81/20 \\
 \hdashline
 10 & 1/20 & 10/20 & 100/20 \\
 \hdashline
 11 & 2/20 & 22/20 & 242/20 \\
 \hdashline
 12 & 3/20 & 36/20 & 432/20 \\
 \hdashline
 13 & 3/20 & 39/20 & 507/20 \\
 \hdashline
 14 & 3/20 & 42/20 & 588/20 \\
 \hdashline
 15 & 3/20 & 45/20 & 675/20 \\
 \hdashline
 16 & 2/20 & 32/20 & 512/20 \\
 \hdashline
 17 & 1/20 & 17/20 & 289/20 \\
 \hdashline
 18 & 1/20 & 18/20 & 324/20 \\
 \hdashline
\end{array}"



Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=\dfrac{270}{20}=13.5=\mu"



The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{3750}{20}-(13.5)^2=5.25= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"

"\sigma_{\bar{X}}=\sqrt{5.25}\approx2.2913"

2.


"\mu_{\bar{X}}=\mu=18.6"

"\sigma_{\bar{X}}=\dfrac{\sigma}{\sqrt{n}}=\dfrac{5.9}{\sqrt{76}}=0.6768"